OpenStudy (ggray):

What is the solution to the equation? (X+8)-14=-15 X=? Will medal

OpenStudy (gibbs):

X=−9

OpenStudy (gibbs):

X+8−14=−15 2 Simplify X+8−14 to X−6 X−6=−15 3 Add 6 to both sides X=−15+6 4 Simplify −15+6 to −9 X=−9

OpenStudy (ggray):

Thx